OpenStudy (kweb2017):

Find the missing term. 7-147 × 798 = 718 × 7-38 × a.7^(-20) b.7^22 c.7^(-29) d.7^58

OpenStudy (anonymous):

is X the missing value or multiplication ?

OpenStudy (kweb2017):

multiplication

OpenStudy (anonymous):

so its 7-147(798)=718(7)-38 ?

OpenStudy (kweb2017):

7^-147*7^98=7^18*7^38*_______

OpenStudy (kweb2017):

@hippo211

OpenStudy (anonymous):

\[7^{-147}(7^{98})=7^{18}(7^{38})\] i this what the problem looks like?

OpenStudy (kweb2017):

yessssss

OpenStudy (kweb2017):

@hippo211

OpenStudy (witty_roger):

i get x = -106

OpenStudy (anonymous):

x is not a value its multiplication

OpenStudy (kweb2017):

i have a multipul choice option..

OpenStudy (witty_roger):

sorry..-105

OpenStudy (witty_roger):

plz check if the question is correct @kweb2017

OpenStudy (kweb2017):

7^-147*7^98=7^18*7^38*_______